OpenStudy (anonymous):

Which equations in point-slope form are equations of the line that pass through the points (4, 5) and (−3, −1)? Choose exactly two answers that are correct.

OpenStudy (anonymous):

A: y - 3 = 7/6 (x - 1) B: y - 5 = 6/7 (x - 4) C: y + 1 = 6/7 (x + 3) D: y - 4 = 7/6 (x - 5)

OpenStudy (apoorvk):

For a line passing through two points say, A\((x_1,y_1)\) and B\((x_2,y_2)\), the equation is: \[\frac{y-y_1}{x-x_1} = \frac{y_1 - y_2}{x_1 - x_2} \] Can you use this for the points given to you? Rearrange in point-slope form after substituting the values.

OpenStudy (anonymous):

how do you do that?

OpenStudy (apoorvk):

You have two points and their coordinates right? A(4,5) and B(-3,-1) (I have named them for simplicity's sake.)

OpenStudy (anonymous):

Ok